Absolute Manipulation​ Manufacturing's (AMM) standards anticipate that there will be 55 pounds of raw material used for every unit of finished goods produced. AMM began the month of MayMay with 4,000 pounds of raw​ material, purchased 27,500 pounds for $19,250 and ended the month with 3,300 pounds on hand. The company produced 5,400 units of finished goods. The company estimates standard costs at $1.10 per pound. The materials price and efficiency variances for the month of May ​were:
Answer #1

Material price variance = (1.1 - 19250/27500)*27500 = $11000 favorable

Material efficiency variance = (5400*55 - (4000+27500-3300)}*1.1 = $1320 unfavorable